# MESTI-MEGANet: Micro-expression Spatio-Temporal Image and Micro-expression Gradient Attention Networks for Micro-expression Recognition

### v1

@@ IV CONCLUSION
In this work, we address the limitations of existing MER methodologies by introducing Micro-expression Spatio-Temporal Image as a novel input modality and Micro-expression Gradient Attention Network as a novel architecture. MESTI effectively encodes micro-movements into a single image, preserving both spatial and temporal features, while MEGANet utilizes a Gradient Attention mechanism to enhance the detection of subtle motion cues.
Our experimental results validate the effectiveness of MESTI by showing that it outperforms all other input modalities, including Apex Frame, Optical Flow, and Dynamic Image, across multiple deep learning networks. Furthermore, replacing the input of previously published MER architectures with MESTI results in significant improvements in recognition accuracy, highlighting its broad applicability. Additionally, MEGANet achieves state-of-the-art performance, particularly when combined with MESTI, confirming its effectiveness in micro-expression analysis. Even when tested with Dynamic Image, MEGANet remains highly competitive, further reinforcing the robustness of its Gradient Attention mechanism. These findings establish MESTI and MEGANet as highly effective solutions for MER, significantly improving recognition accuracy. Future work could explore refining MESTI for real-time applications, integrating additional attention mechanisms, or leveraging larger-scale datasets to further advance micro-expression recognition systems.

### v2

@@ IV CONCLUSION
In this work, we address the limitations of existing MER methodologies by introducing ME Spatio-Temporal Image as a novel input modality and ME Gradient Attention Network as a novel architecture. MESTI effectively encodes micro-movements into a single image, preserving both spatial and temporal features, while MEGANet utilizes a Gradient Attention mechanism to enhance the detection of subtle motion cues.
Our experimental results validate the effectiveness of MESTI by showing that it outperforms all other input modalities, including Apex Frame, Optical Flow, and Dynamic Image, across multiple deep learning networks. Furthermore, replacing the input of previously published MER architectures with MESTI results in significant improvements in recognition accuracy, highlighting its broad applicability. Additionally, MEGANet achieves state-of-the-art performance, particularly when combined with MESTI, confirming its effectiveness in ME analysis. These findings establish MESTI and MEGANet as highly effective solutions for MER, significantly improving recognition accuracy. Future work could explore refining MESTI for real-time applications, integrating additional attention mechanisms, or leveraging larger-scale datasets to further advance ME recognition systems.